[figure: 9.7 cm]

Transport and magnetic properties of Nd_{1−x}A_{x}MnO_{3}/YBa_{2}Cu_{3}O_{7}

heterostructures

I. Komissarov, E. Dynowska, M. Sawicki, A. Szewczyk, P. Przysłupski Institute of Physics, PAS, Al. Lotników 32/46, Warsaw 02-668

Ferromagnetic (F) and superconducting (S) phase are antogonistic to each other. However, a coexistence ferromagnetic and superconducting phase was observed in Nd_{1−x}A_{x}MnO_{3}/YBa_{2}Cu_{3}O_{7} (NSMO/YBCO) superlatices, with modulation length less then 10 nm. In the present study we report on the growth, structure and magnetic characterization of (NSMO/YBCO) superlattices. The comporative studies for two series of NSMO/YBCO superlattices have been performed i.e. for NSMO with x=0.19 and 0.33 respectiviely. For the doping level x=0.19 the bulk NSMO is a ferromagnetic insulator (FMI) with Curie temperature of about 100 K. Whereas for x=0.33 NSMO is a ferromagnetic metal, (FMM), with Curie temperature of about 200 K. Two seris of NSMO/YBCO superlattices for both doping level were were deposited with fixed NSMO layer equal to 4.8 nm thickness and varying YBCO thickness from 3.6 nm up to 8.4 nm. X ray diffraction spectrum show the presence of superstructure reflected satelittes indicating a good quality of the S/F interfaces in multilayered heterostructures. Neverthless, trasport measurements show transition to superconducting state for multilayers with the YBCO layer thickness above 3.6 nm. The multilayers show in-plane magnetic anisotropy above and below superconducting transition. Discussion of the experimental results for the two ferromagnetic subsystem in NSMO/YBCO heterostructure on their superconducting properties will be presented. Supported by State Committee for Scientific Research,(KBN), grant No 5P03B6220.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : I. Komissarov

Address for correspondence : Institute of Physics Polish Academy of Sciences Al. Lotników 32/46, Warsaw 02-668

Email address : komis@ifpan.edu.pl